OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 less than or greater to x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

-3x+4<=x+8 -3x-x+4<=x-x+8 -4x+4<=8 -4x<=4 x>=-1

OpenStudy (anonymous):

Thank you! :)